Amy paid $89.65 for a pair of running shoes during a 15%-off sale. What was the regular price?
Answer by mananth(16946)

15 % discount means she pays 85 if the regular price is 100
...
100/85 = x/89.65
100*89.65/85=105.47
regular price = $105.47